OpenStudy (anonymous):

When adding the fractions, do I have this right? 3/p+3 + 4/p = 7/p+3 ???

OpenStudy (anonymous):

not exactly

OpenStudy (anonymous):

p^2 + 3??

OpenStudy (anonymous):

You have to have the same denominator before you add them

OpenStudy (dumbcow):

to add these fractions, make common denominator p*(p+3)

OpenStudy (anonymous):

10/p?

OpenStudy (anonymous):

I meant 7/p? gosh, lol. shooting in the dark!

jimthompson5910 (jim_thompson5910):

\[\frac{3}{p+3}+\frac{4}{p}\] \[\frac{3p}{p(p+3)}+\frac{4}{p}\] \[\frac{3p}{p(p+3)}+\frac{4(p+3)}{p(p+3)}\] \[\frac{3p}{p(p+3)}+\frac{4p+12}{p(p+3)}\] \[\frac{3p+4p+12}{p(p+3)}\] \[\frac{7p+12}{p(p+3)}\] \[\frac{7p+12}{p^2+3p}\] So \[\frac{3}{p+3}+\frac{4}{p}=\frac{7p+12}{p^2+3p}\]